Session: Feb 2012 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Bo Xia at Fri Mar 02 12:00:12 2012) Success:

Svetloe (T.Andreeva at Fri Mar 02 12:18:35 2012) Success:
Very bad weather: wet snow

Torun (Kaz Borkowski at Fri Mar 02 12:30:24 2012) Minor failures:
BBC7 unlocked during the entire experiment. No other known problems.

Hartebeesthoek (Jonathan Quick at Fri Mar 02 13:07:37 2012) Success:
No known problems apart from the usual RFI

Medicina (Monari at Fri Mar 02 13:46:53 2012) Success:

Westerbork (Geert Kuper at Fri Mar 02 17:01:11 2012) Success:
No known problems

Onsala (Michael Lindqvist at Mon Mar 05 11:01:14 2012) Success:
Problem with the noise diode first scans, otherwise no known problems. Finally fringes with the DBBC again. Input level to low previously??

Effelsberg (Uwe Bach at Tue Mar 06 09:03:00 2012) Success:

Urumqi (Hua Zhang at Thu Mar 08 12:33:51 2012) Success:

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:15:59 2012) Success:
No known problems.